OpenStudy (anonymous):

A cable installer charges $42.50 per hour plus a $45.00 service charge. Your father's firm hires him to hook up his company's Internet service. Find the total charges if it takes the cable installer 5 hours to complete the task.

OpenStudy (anonymous):

$42.50* 5= 212.5 212.5+44=?

OpenStudy (anonymous):

42.5(5)+45=

OpenStudy (anonymous):

Thank you i got it (:

OpenStudy (anonymous):

Your welcome :)